The number of LEDs required to set something on fire is given by

\frac{(A − T)mc}{tP fa}

where

• A is the autoignition point of the fuel

• T is the ambient temperature

• m is the mass of the fuel

• c is the specific heat capacity of the fuel

• t is the time available for ignition

• P is the power emitted by one LED

• f is the fraction of light energy focused onto the fuel by the lens system • a is the fraction of light energy absorbed by the fuel. This formula assumes (among other things) that a does not change appre-

ciably with the temperature of the fuel, and that the fuel does not lose heat to its environment.